# LLMs are Also Effective Embedding Models: An In-depth Overview

### v1

@@ 8 Open Problems
Embedding Quality Across Different Tasks. While LLMs have shown impressive performance in generating embeddings for specific tasks, such as semantic similarity or text classification, their effectiveness across a broader range of tasks remains uncertain. Tasks like clustering, reranking, and summary require embeddings to capture different kinds of relationships between data points, and it is not always clear if embeddings generated by LLMs can consistently perform well in these varied contexts. Moreover, different tasks may prioritize different aspects of the embedding space, such as local versus global structure or task-specific nuances (Cai et al., 2022), which might not be fully captured by general-purpose LLM embeddings. This raises the open question of how LLM embeddings can be optimized or adapted to maintain high-quality representations across diverse tasks (Li et al., 2024d), and whether task-specific tuning or hybrid approaches might be necessary to improve overall effectiveness in more complex and specialized applications.
Efficiency vs. Accuracy Trade-offs. While LLM-generated embeddings can achieve better accuracy in capturing semantic relationships, they often come at a significant computational cost due to the large model sizes and high-dimensional representations (e.g., 4096 for most recent LLMs). This creates a critical challenge in balancing efficiency and accuracy. For tasks that require real-time processing or deployment in resource-constrained environments, the time and memory demands of using LLMs can be prohibitive. Reducing the computational burden without sacrificing the quality of embeddings is an open problem. Techniques such as model distillation (Lee et al., 2024b), pruning, or dimensionality reduction have been proposed, but each comes with trade-offs in terms of performance loss or reduced representational power. Moreover, it remains unclear to what extent the high-dimensional nature of LLM embeddings contributes to their effectiveness, and whether lower-dimensional embeddings could perform comparably in specific tasks. Thus, research is needed to explore strategies that make LLM embeddings more efficient while preserving their robustness, especially in applications where both accuracy and real-time performance are critical. Additionally, exploring the embedding performance of relatively smaller models, such as Phi3 (Abdin et al., 2024) and MiniCPM (Hu et al., 2024), could offer valuable insights into achieving a balance between computational efficiency and embedding quality.
@@
Adapting LLM Embeddings for Low-Resource Domains. One of the major challenges in using LLM-based embeddings is their adaptability to low-resource domains, where training data is sparse, or under-representative of the target domain. While LLMs have been pre-trained on large, diverse datasets, these models often struggle when applied to specialized domains, such as medical, legal, or technical fields where pair data is limited (Li et al., 2024c; Sivarajkumar et al., 2024). The embeddings generated in these settings may fail to capture the necessary domain-specific nuances, leading to limited performance on downstream tasks like question answering or retrieval. Fine-tuning LLMs for such low-resource domains requires domain-specific data that may not always be available. Additionally, overfitting to small datasets is a concern, which raises the question of how to adapt LLM embeddings to low-resource settings without sacrificing generalization.
@@
How to effective training for embedding models? LLMs are typically pre-trained on large text corpora for general language modeling tasks, but adapting them specifically for generating high-quality embeddings often requires resource-intensive contrastive training. This poses a significant challenge, particularly in scenarios where only limited data is available. An intriguing open problem is how to effectively tune LLMs for embedding tasks using small-scale datasets. For example, LIMA demonstrates that only about 1,000 high-quality, human-curated samples are sufficient for aligning LLMs in general supervised fine-tuning (Zhou et al., 2023). This suggests that efficient fine-tuning with a smaller amount of data might be feasible for embedding tasks as well. Additionally, exploring whether there are more effective training methods or strategies tailored specifically for tuning LLMs as embedding models could provide valuable insights.
@@ 9 Conclusion
[…]

### v2

@@ 8. Open Problems
Embedding Quality Across Different Tasks. While LLMs have shown impressive performance in generating embeddings for specific tasks, such as semantic similarity or text classification, their effectiveness across a broader range of tasks remains uncertain. Tasks like clustering, reranking, and summary require embeddings to capture different kinds of relationships between data points, and it is not always clear if embeddings generated by LLMs can consistently perform well in these varied contexts. Moreover, different tasks may prioritize different aspects of the embedding space, such as local versus global structure or task-specific nuances (Cai et al., 2022), which might not be fully captured by general-purpose LLM embeddings. This raises the open question of how LLM embeddings can be optimized or adapted to maintain high-quality representations across diverse tasks (Li et al., 2024b), and whether task-specific tuning or hybrid approaches might be necessary to improve overall effectiveness in more complex and specialized applications.
Efficiency vs. Accuracy Trade-offs. While LLM-generated embeddings can achieve better accuracy in capturing semantic relationships, they often come at a significant computational cost due to the large model sizes and high-dimensional representations (e.g., 4096 for most recent LLMs). This creates a critical challenge in balancing efficiency and accuracy. For tasks that require real-time processing or deployment in resource-constrained environments, the time and memory demands of using LLMs can be prohibitive. Reducing the computational burden without sacrificing the quality of embeddings is an open problem. Techniques such as model distillation (Lee et al., 2024a), pruning, or dimensionality reduction have been proposed, but each comes with trade-offs in terms of performance loss or reduced representational power. Moreover, it remains unclear to what extent the high-dimensional nature of LLM embeddings contributes to their effectiveness, and whether lower-dimensional embeddings could perform comparably in specific tasks. Thus, research is needed to explore strategies that make LLM embeddings more efficient while preserving their robustness, especially in applications where both accuracy and real-time performance are critical. Additionally, exploring the embedding performance of relatively smaller models, such as Phi3 (Abdin et al., 2024) and MiniCPM (Hu et al., 2024), could offer valuable insights into achieving a balance between computational efficiency and embedding quality.
@@
Reasoning Gaps in Embedding-Based Retrieval. Despite progress in retrieval modeling, current systems still struggle to perform robustly on reasoning-intensive tasks. A fundamental issue is the lack of reasoning capability in the first-stage retrievers, which are responsible for high-recall candidate generation. These retrievers, typically based on dense or hybrid embeddings, often fail to capture implicit connections, logical dependencies, or multi-step inference chains required for identifying relevant documents. As a result, important evidence is missed early on, limiting the performance ceiling of downstream rerankers or generators. This challenge is especially prominent in benchmarks like BRIGHT and ImpliRet, where retrieval requires understanding implicit facts or combining dispersed clues. The absence of fine-grained supervision for reasoning types during training further compounds the issue, as retrievers are usually optimized with weak relevance signals that do not reflect the nature of the reasoning required. Moreover, standard IR metrics such as recall or nDCG are not designed to evaluate reasoning quality, making progress difficult to measure. Addressing these challenges will require retrieval models that are not only semantically aware but also reasoning-aware—capable of aligning with the latent inference patterns embedded in user queries.
@@
Adapting LLM Embeddings for Low-Resource Domains. One of the major challenges in using LLM-based embeddings is their adaptability to low-resource domains, where training data is sparse, or under-representative of the target domain. While LLMs have been pre-trained on large, diverse datasets, these models often struggle when applied to specialized domains, such as medical, legal, or technical fields where pair data is limited (Li et al., 2024f; Sivarajkumar et al., 2024). The embeddings generated in these settings may fail to capture the necessary domain-specific nuances, leading to limited performance on downstream tasks like question answering or retrieval. Fine-tuning LLMs for such low-resource domains requires domain-specific data that may not always be available. Additionally, overfitting to small datasets is a concern, which raises the question of how to adapt LLM embeddings to low-resource settings without sacrificing generalization.
@@
How to effective training for embedding models? LLMs are typically pre-trained on large text corpora for general language modeling tasks, but adapting them specifically for generating high-quality embeddings often requires resource-intensive contrastive training. Recent models increasingly rely on large-scale paired data for contrastive fine-tuning; for instance, Qwen3-Emb is trained on 150 million paired examples (Zhang et al., [n. d.]). This poses a significant challenge, particularly in scenarios where only limited data is available. An intriguing open problem is how to effectively tune LLMs for embedding tasks using small-scale datasets. For example, LIMA demonstrates that only about 1,000 high-quality, human-curated samples are sufficient for aligning LLMs in general supervised fine-tuning (Zhou et al., 2023). This suggests that efficient fine-tuning with a smaller amount of data might be feasible for embedding tasks as well. Additionally, exploring whether there are more effective training methods or strategies tailored specifically for tuning LLMs as embedding models could provide valuable insights.
@@ 9. Conclusion
Acknowledgements.
We thank the editors and the anonymous reviewers for their constructive comments and advice.